OpenStudy (anonymous):

I am back! I am stuck on the same problem, but at least I am further along:)

OpenStudy (anonymous):

\[x=+4\pm \sqrt{16+8} all of this is divided by 2\]

OpenStudy (anonymous):

What is 16 + 8?

OpenStudy (anonymous):

See, that is what we were unsure of; do we add or take the square root of both numbers?

OpenStudy (anonymous):

For instance, would it be 24 or do you take the square root of the 16 and the 8?

OpenStudy (anonymous):

You have to add first, then take the square root. \[\sqrt{a + b} \ne \sqrt{a} + \sqrt{b}\]

OpenStudy (anonymous):

But then you have to break it down again? What do you do with the number outside?

OpenStudy (anonymous):

Well what is \(\sqrt{24}\)

OpenStudy (anonymous):

4*6

OpenStudy (anonymous):

Or 2\[\sqrt{6}\]

OpenStudy (anonymous):

Right. \(2\sqrt{6}\). So you have: \[{4 \pm 2\sqrt{6} \over 2}\]

OpenStudy (anonymous):

Ok, can you then reduce it down?

OpenStudy (anonymous):

Of course. You can divide each of the terms by 2. \[{4 \pm 2\sqrt{6} \over 2} = {4 \over 2} \pm {2\sqrt{6} \over 2} = 2 \pm \sqrt{6}\]

OpenStudy (anonymous):

Yea, we got that far:) the problem we have is getting it into decimal form.

OpenStudy (anonymous):

Why bother?

OpenStudy (anonymous):

Oh, we just figured it out! Thank you thank you thank you!!

OpenStudy (anonymous):

I mean if you want you can just figure out what the decimal form of \(\sqrt{6} \) is then add and subtract that from 2 and you have your two solutions.. But the original way looks lots neater.

OpenStudy (anonymous):

Most certainly. The answer according to the professor is in decimal form which is why we were wondering. Thank you thank yyou thank you!! We just figured it out!

OpenStudy (anonymous):

Nicely done! =)

OpenStudy (anonymous):

YOu are great! Thank you so much! GOD bless<><

OpenStudy (anonymous):

I'm technically Buddhist, but thanks anway ;)

OpenStudy (anonymous):

Well thank you, GOD blessed us through you.

OpenStudy (anonymous):

Fair enough =)

OpenStudy (anonymous):

God loves and appreciates you as do we:)

OpenStudy (anonymous):

Happy to help.